OpenStudy (anonymous):

when rolling 2 dice what is the probability of getting a total of 8 or 12?

OpenStudy (anonymous):

P(A or B) = P(A) + P(B) So what's the probability of getting an 8?

OpenStudy (anonymous):

Or getting a total of 8 rather.

OpenStudy (blockcolder):

There are 36 possibilities in total. The probability of getting an 8 is 5/36 ((2,6)(3,5)(4,4)(5,3)(6,2)). The probability of getting a 12 is 1/36. (6,6) The probability of getting either is P(A)+P(B)-P(A and B)=5/36+1/36-0(since you can't get an 8 and a 12 at the same time)=6/36 or 1/6. :D

OpenStudy (anonymous):

Yeah, I dunno why I bother either.